how to get the answer of the following?
1.) The perimeter of an isosceles triangle is 36 cm. The total length of the two congruent legs is 12 centimeters more than one half the length of the base. Find the lengths of the three sides.

Answer by mananth(16946)

let the base = x
let congruent side be y
2y+x=36 .....................1
2y= x/2 +12
multiply by 2
4y=x+24
4y-x=24.......................2
add eq 1 & eq 2
6y=36+24
6y=60
divide by 6
y= 60/6
y=10 conguent side
plug value of y in equation 1
2*10+x=36
20+x=36
x=36-20
x=16 the base
CHECK
16+10+10 = 36
.....